# Thermal Modelling of Battery Cells for Optimal Tab and Surface Cooling Control

### v1

@@ IV-D Results of comparison against the TEC model
To investigate the efficacy and applicability of the CSG model validated in Section IV-C, we compare it with the second type of benchmark defined in Section IV-A, i.e., the TEC model. In the TEC model, the average temperature is derived as $T^{\text{TEC}}_{\text{mean}}=(T_{s}+T_{c})/2$, and the radial gradient is defined by $\delta T^{r}_{max}=(T_{c}-T_{s})/(R_{\text{out}}-R_{\text{in}})$. To possibly fully reveal their predictive capability, particularly under highly dynamic operating conditions, we introduce a pulsed power load profile to generate the heat $q(t)$, as shown in Fig. 3a. This heat trajectory is characterised by high heat generation rates but a relatively short duration. The comparative results between the FEM, CSG, and TEC models under surface cooling are presented in Fig. 3.
@@
It is found that the CSG model of $O$ = 1 is capable of closely following the true trajectories of the mean temperature, max temperature, and max gradient in the entire test cycle. By contrast, in general, the TEC model with an order of 2 significantly underestimates all three considered indicators, resulting in a very large error. To observe more closely, the model’s error for $T_{\max}$ is much larger than that for $T_{\text{mean}}$, and it increases when there are abrupt changes in the heat generation. This means that compared to our developed model, the TEC model is unreliable for predicting the local thermal behaviour and may be limited to applications with only small heat generation in the battery cell. Particularly, this model cannot be used for battery safety management where the maximum temperature must be accurately estimated and constrained, and for thermal uniformity control where the information on the maximum gradients is critical.
@@
Computational efficiency is of great importance for model-based applications, particularly real-time thermal optimization and control of large-scale battery packs. With this in mind, we examine the computational times of the developed model with various orders and compare them with that of the TEC model. To have a fair comparison, we run each model ten times and compute their average simulation times in milliseconds (ms), with the results listed in Table I. Clearly, the CSG model with an order $O$=1 is more efficient than its TEC counterpart, leading to a reduction of 28.7% in the required time. The model with higher orders takes longer to simulate, but its time consumption remains within the same order of magnitude as that of the TEC model.
@@
The above results indicate that the CSG model with $O$=1 offers distinct advantages in both accuracy and computational efficiency. Therefore, it can readily replace the TEC model in existing BMSs to improve battery thermal performance in the real world. In addition, it may be computationally feasible to use a high-order CSG model, e.g., the order of 4, within a BMS, leading to significantly advanced management of battery safety and lifetime.
@@
TABLE I: Average computational times of CSG model of different orders compared with the TEC model.
Model type TEC $O$ = 1 $O$ = 4 $O$ = 9 $O$ = 16 $O$ = 25
Time [ms] 30.7 21.9 43.1 51.7 65.2 74.6
@@ V Model-based applications
The developed models for cylindrical and pouch battery cells, featuring high fidelity and efficiency, can be effectively applied to various model-based applications. These include rapid analysis of different cooling scenarios, real-time closed-loop temperature control, and thermal optimization of cell design, as demonstrated in this section.
@@ V-A Model-based analysis for different cooling scenarios
For a given battery system and usage profiles, selecting the most appropriate cooling scenario can make a big difference to all its key properties. We analyze thermal performance merits, $T_{\text{mean}}$, $T_{\max}$, $\delta T^{r}_{\max}$, $\delta T^{z}_{\max}$, and $\Delta T=T_{\max}-T_{\min}$, under the five cooling scenarios defined in Section IV-B, with the results for cylindrical cells presented in Table II.
In terms of $T_{\text{mean}}$, bTC gives rise to the highest value at 22.77 C, followed by SC with the second-highest value. This is because the average temperature is primarily influenced by the cooling area, and both bTC and SC have only one side exposed to cooling. With the largest cooling area, aTSC demonstrates a high capability of heat removal, leading to the lowest values of $T_{\text{mean}}$ and $T_{\max}$ among all cases. However, aTSC does not achieve the lowest thermal gradient. In contrast, btTC provides the lowest $\delta T^{r}_{\max}$ at the local level and the lowest $\Delta T$ across the entire cell. This can be attributed to symmetric cooling from the cell’s top and bottom sides, and that all electrodes get access to cooling at the same time under btTC because of the rolled electrode layer structure. On the contrary, in the case of SC, the outer layers get cooled before the internal layers, leading to inhomogeneous cooling and consequently the highest gradient of 32.1 C/m.
@@
TABLE II: Thermal performance merits under the five scenarios.
@@ V-B Closed-loop temperature control
Since $T_{\text{mean}}$ is not measurable during real-world battery operation, it is imperative to incorporate a state estimator within the designed control strategy. In this preliminary study, we design an open-loop estimator that mirrors the model (39)–(40) for cylindrical battery cells. Beyond its high accuracy and efficiency, this thermal model enables precise control of cooling applied to any side of the cell or combinations thereof, which is a capability not offered by existing models such as those in [15, 16]. This flexibility makes our model particularly suitable for battery temperature estimation and control in sophisticated cooling scenarios.
@@
Figure 5: Results of battery thermal control. (a) and (b) show the results obtained in aTSC and the 1C case of WLTP. (c) and (d) compare the mean values of the radial and axial thermal gradients, respectively, for different cooling scenarios and WLTP cases.
For case studies, the original WLTP profile is scaled so that the maximum current becomes 1, 2, 3, and 4C, respectively. The control results are given in Fig. 5. To save space, in Fig. 5a and b, only the trajectories obtained in aTSC and the 1C case are depicted.
Based on the results in Fig. 5c and d, one can readily find the answer to the question raised. When the average cell temperature is maintained at 20 C, btTC best balances the thermal gradients along the $r$ and $z$ directions and produces the most homogeneous thermal distribution across the cylindrical cell. Specifically, in btTC, the average thermal gradient along the $r$ direction, denoted by $\delta T^{r}_{\text{mean}}$, is substantially smaller than those observed in all other cooling scenarios. Regarding the average thermal gradient along the $z$ direction, btTC is comparable to the best-performing scenario, i.e., SC.
In terms of the control’s behaviour, as demonstrated in Fig. 5a and b, substantial control efforts are initially required to attenuate the large tracking errors. Once $T_{\text{mean}}$ approaches the set-point, the magnitude of $u_{t}$, $u_{b}$, and $u_{s}$ decreases significantly. More sophisticated control strategies will be explored and implemented in future work to systematically enhance the control performance, in which multiple control objectives and potential constraints will be explicitly addressed. Within the control strategy, the new models developed in Section III can serve as efficient tools for accurately simulating and predicting battery thermal states.
@@ V-C Model-based evaluation and enhancement for cell design
The results show that under the cooling scenarios bTC, bTSC, btTC, and aTSC, a higher value of $L/R_{\text{out}}$ typically leads to a more pronounced temperature rise and greater thermal gradients across the battery cell. This means that the cell’s thermal behaviour is highly sensitive to its dimensions. Under these four scenarios, short and bulky battery cells are more favourable compared to their tall and slim alternatives. Specifically, the reduced mean temperature can help mitigate battery ageing mechanisms such as solid-electrode interface (SEI) growth, while the more uniform thermal distribution promotes more homogeneous usage of active electrode materials, both of which contribute to a longer cell lifespan. These findings can be used to evaluate existing cell designs. For battery types of 18650, 26650, 21700, and 4680 prevalent in today’s market, their $L/R_{\text{out}}$ values are 7.22, 5.42, 6.67, and 3.48, respectively. It is evident that the 4680 cell type offers the best thermal performance across the four scenarios involving at least one tab in the cooling process.
@@
Figure 6: Evaluation of the thermal performance of cylindrical cells with varying dimensions under five cooling scenarios. (a) Evolution of the mean temperature $T_{\text{mean}}$. (b) Evolution of the max thermal gradients in the radial direction $\delta T^{r}_{\text{max}}$. (c) Evolution of the max thermal gradients in the axial direction $\delta T^{z}_{\text{max}}$.

### v2

@@ IV-D Results of comparison against the TEC model
To investigate the efficacy and applicability of the CSG model validated in Section IV-C, we compare it with the second type of benchmark defined in Section IV-A, i.e., the TEC model. In the TEC model, the average temperature is derived as $T^{\text{TEC}}_{\text{mean}}=(T_{s}+T_{c})/2$, and the radial gradient is defined by $\delta T^{r}_{max}=(T_{c}-T_{s})/(R_{\text{out}}-R_{\text{in}})$. To possibly fully reveal their predictive capability, particularly under highly dynamic operating conditions, we introduce a pulsed power load profile to generate the heat $q(t)$, as shown in Fig. 3(a). This heat trajectory is characterised by high heat generation rates but a relatively short duration. For completeness, the TEC model was also included in the WLTP validation scenario shown in Fig. 2(b) and (c). The comparative results between the FEM, CSG, and TEC models under the pulsed power profile with surface cooling are presented in Fig. 3.
@@
It is found that the CSG model of $O$ = 1 is capable of closely following the true trajectories of the mean temperature, max temperature, and max gradient in the entire test cycle. By contrast, in general, the TEC model with an order of 2 significantly underestimates all three considered indicators, resulting in temperature errors of several degrees. To observe more closely, the model’s error for $T_{\max}$ is much larger than that for $T_{\text{mean}}$, and it increases when there are abrupt changes in the heat generation. This observation is consistent with the comparison under the WLTP profile in Fig. 2(b) and (c), where the TEC model similarly underestimates the core and surface temperatures. This means that compared to our developed model, the TEC model is unreliable for predicting the local thermal behaviour and may be limited to applications with only low heat generation in the battery cell. Particularly, this model cannot be used for battery safety management, where the maximum temperature must be accurately estimated and constrained, and for thermal uniformity control, where the information on the maximum gradients is critical.
@@
Computational efficiency is of great importance for model-based applications, particularly real-time thermal optimization and control of large-scale battery packs. With this in mind, we examine the computational times of the developed model with various orders and compare them with that of the TEC and FEM models. To have a fair comparison, we run each model ten times and compute their average simulation times in milliseconds (ms), with the results listed in Table II. Clearly, the CSG model with an order $O$=1 is more efficient than its TEC counterpart, leading to a reduction of 28.7% in the required time. The model with higher orders takes longer to simulate, but its time consumption remains within the same order of magnitude as that of the TEC model. As expected, the FEM model takes a longer computational time, about 8s to compute the solution, making it impractical for real-time control applications.
@@
The above results indicate that the CSG model with $O$=1 offers distinct advantages in both accuracy and computational efficiency. Therefore, it can readily replace the TEC model in existing BMSs to improve battery thermal performance in the real world. In addition, it may be computationally feasible to use a high-order CSG model, e.g., the order of 4, within a BMS, allowing a significantly more advanced management of battery safety and lifetime.
@@
TABLE II: Average computational times of CSG model of different orders compared with the TEC model.
Model type FEM TEC $O$=1 $O$=4 $O$=9 $O$=16 $O$=25
Time [ms] 8000 307 219 431 517 652 746
@@ V Model-based applications
The developed models for cylindrical and pouch battery cells, featuring high fidelity and efficiency, can be effectively applied to various model-based applications. These include rapid analysis of different cooling scenarios, real-time closed-loop temperature control, and thermal optimization of cell design, as demonstrated in this section. Here, a model order $O=9$ is selected and simulations are conducted under the mWLTP profile on the battery cell and cooling scenarios specified in Section IV-B.
@@ V-A Model-based analysis for different cooling scenarios
For a given battery system and usage profiles, selecting the most appropriate cooling scenario can make a big difference to all its key properties. We analyze thermal performance merits, $T_{\text{mean}}$, $T_{\max}$, $\delta T^{r}_{\max}$, $\delta T^{z}_{\max}$, and $\Delta T=T_{\max}-T_{\min}$, under the five cooling scenarios defined in Section IV-B, with the results for cylindrical cells presented in Table III.
In terms of $T_{\text{mean}}$, bTC gives rise to the highest value at 22.77 C, followed by SC with the second-highest value. This is because the average temperature is primarily influenced by the cooling area, and both bTC and SC have only one side exposed to cooling. With the largest cooling area, aTSC demonstrates a high capability of heat removal, leading to the lowest values of $T_{\text{mean}}$ and $T_{\max}$ among all cases. However, aTSC does not achieve the lowest thermal gradient. In contrast, btTC provides the lowest $\delta T^{r}_{\max}$ at the local level and the lowest $\Delta T$ across the entire cell. This can be attributed to symmetric cooling from the cell’s top and bottom sides, and that all electrodes get access to cooling at the same time under btTC because of the rolled electrode layer structure. On the contrary, in the case of SC, the outer layers get cooled before the internal layers, leading to inhomogeneous cooling and consequently the highest gradient of 32.1 C/mm.
@@
TABLE III: Thermal performance merits under the five scenarios.
@@ V-B Closed-loop temperature control
Since $T_{\text{mean}}$ is not measurable during real-world battery operation, it is imperative to incorporate a state estimator within the designed control strategy. In this preliminary study, we design an open-loop estimator that mirrors the model (38)–(39) for cylindrical battery cells. However, the state-space model allows any standard state-estimator such as the Kalman filter. Beyond its high accuracy and efficiency, this thermal model enables precise control of cooling applied to any side of the cell or combinations thereof, which is a capability not offered by existing models such as those in [17, 18]. This flexibility makes our model particularly suitable for battery temperature estimation and control in sophisticated cooling scenarios.
@@
Figure 5: Results of battery thermal control of the cylindrical cell. (a) and (b) show the results obtained in aTSC and the 1C case of WLTP. (c) and (d) compare the mean values of the radial and axial thermal gradients, respectively, for different cooling scenarios and WLTP cases.
For case studies, the original WLTP profile is scaled so that the maximum current becomes 1, 2, 3, and 4C, respectively. The control results are given in Fig. 5. To save space, in Fig. 5(a) and (b), only the trajectories obtained in aTSC and the 1C case are depicted.
Based on the results in Fig. 5(c) and (d), one can readily find the answer to the question raised. When the average cell temperature is maintained at 20 C, btTC best balances the thermal gradients along the $r$- and $z$-directions and produces the most homogeneous thermal distribution across the cylindrical cell. Specifically, in btTC, the average thermal gradient along the $r$-direction, denoted by $\delta T^{r}_{\text{mean}}$, is substantially smaller than those observed in all other cooling scenarios. Regarding the average thermal gradient along the $z$-direction, btTC is comparable to the best-performing scenario, i.e., SC.
In terms of the control’s behaviour, as demonstrated in Fig. 5(a) and (b), substantial control efforts are initially required to attenuate the large tracking errors. Once $T_{\text{mean}}$ approaches the set-point, the magnitude of $u_{t}$, $u_{b}$, and $u_{s}$ decreases significantly. More sophisticated control strategies will be explored and implemented in future work to systematically enhance the control performance, in which multiple control objectives and potential constraints will be explicitly addressed. Within the control strategy, the new models developed in Section III can serve as efficient tools for accurately simulating and predicting battery thermal states.
@@ V-C Model-based evaluation and enhancement for cell design
The results show that under the cooling scenarios bTC, bTSC, btTC, and aTSC, a higher value of $L/R_{\text{out}}$ typically leads to a more pronounced temperature rise and greater thermal gradients across the battery cell. This means that the cell’s thermal behaviour is highly sensitive to its dimensions. Under these four scenarios, short and bulky battery cells are more favourable compared to their tall and slim alternatives. Specifically, the reduced mean temperature can help mitigate battery ageing mechanisms such as solid-electrode interface (SEI) growth, while the more uniform thermal distribution promotes more homogeneous usage of active electrode materials, both of which contribute to a longer cell lifespan [3, 36, 37]. These findings can be used to evaluate existing cell designs. For battery types of 18650, 26650, 21700, and 4680 prevalent in today’s market, their $L/R_{\text{out}}$ values are 7.22, 5.42, 6.67, and 3.48, respectively. It is evident that the 4680 cell type should offer the best thermal performance across the four scenarios involving at least one tab in the cooling process.
@@
Figure 6: Evaluation of the thermal performance of cylindrical cells with varying dimensions under five cooling scenarios. (a) Evolution of the mean temperature $T_{\text{mean}}$. (b) Evolution of the maximal thermal gradients in the radial direction $\delta T^{r}_{\text{max}}$. (c) Evolution of the maximal thermal gradients in the axial direction $\delta T^{z}_{\text{max}}$.